The surface of water in a water tank on the top of a house is 4m above the tap level . Find the pressure of water at the tap when the tap is clised.

Grade:11

1 Answers

Xyz
6 years ago
Density of water =1000 kg/M3Height between surface of water and closed tap = 4mAcceleration due to gravity =9.8m/S2Pressure =Density×height×g=4×1000×9.8=39200